# A Survey of LLM-based Automated Program Repair: Taxonomies, Design Paradigms, and Applications

### v1

@@ 5.4. Discussion and Outlook
Future research in agentic repair should emphasize cost-aware planning that caches intermediate results, prunes unproductive branches, and calls judgment modules only when their expected benefit justifies the extra computational resources, thereby reducing latency. Robust context management is also essential: agents must retain only the most relevant code fragments, tests, and analysis artifacts so that large repositories remain within the model’s context window without losing critical detail. Reliability will improve when lightweight static checks and policy guards intercept infinite loops or unsafe edits before they reach version control, satisfying continuous integration requirements. Finally, a tiered workflow that applies inexpensive heuristics to straightforward defects and escalates more complex cases to a fully autonomous agent can strike a balance between efficiency and autonomy. Sustained attention to these directions can elevate current prototypes into dependable tools for everyday software engineering practice.
Table 3. Detailed Information of LLM-based APR Systems
@@
Fine-Tuning Full Fine-Tuning Huang (2023) Recoder dataset CodeT5 Defects4J v1.2/2.0 pass@5 82/107 Perfect Fault Localization
Jiang (2023) Recoder dataset Incoder 7B Defects4J v1.2/2.0 pass@10 31/138 Perfect Fault Localization
@@
Fine-Tuning PEFT RepairLlama Megadiff CodeLlama 7B HumanEval-Java pass@10 67.28% Perfect Fault Localization
@@
Luo (2024) TutorCode CodeLlama 13B HumanEval-Java pass@10 76.88%
Li (2024) APR Instruction DeepSeek-Coder 7B HumanEval-Java pass@10 68.10% Perfect Fault Localization
Ruiz (2025) Recoder dataset DeepSeek-Coder 7B HumanEval-Java pass@10 78.53% Perfect Fault Localization
Fine-Tuning Knowledge Distilling KNOD GitHub / Defects4J v1.2/2.0 pass@1000 122/837 Perfect Fault Localization
DistiLRR MBXP CodeLlama-7B HumanEval-Perl pass@1 42.10% train/test split
@@
Fine-Tuning RLFT RePair CodeNet4Repair StarCoder-16B CodeNet4Repair pass@5 65.66% train/test split
@@
Smart-LLaMA-DPO built trainset Llama3.1 8B test split F1 83.4 train/test split
Fine-Tuning Context-Enriched TraceFixer CodeNet CodeT5-small CodeNet pass@10 87%
InferFix InferredBugs CodeX 12B InferredBugs pass@1 76.80% train/test split
@@
Prompting Zero-Shot AlphaRepair CodeBERT Defects4J v1.2/2.0 top@k (5h) 74/473 Perfect Fault Localization
Prenner (2022) CodeX-175B Quixbugs pass@1 37/80 Perfect Fault Localization
Fan (2023) CodeX-175B LMDefects pass@5 46/113
Tian (2023) GPT-3.5 Turbo LeetCode pass@5 105/120
Prompting Few-Shot Xia (2023) CodeX 12B Defects4J v1.2 pass@600 99/255 Perfect Fault Localization
Gao (2023) CodeX-175B B2F small full match 32.25%
Ahmed (2023) CodeX-175B MODIT pass@30 31.80%
@@
Prompting + RAG T-RAP CodeT5 Big-Fixes pass@10 25.80% Perfect Fault Localization
@@
Prompting + AAG D4C GPT-4 Defects4J v1.2/2.0 pass@10 180/437 single-function bugs
@@
Procedural Test-in-the-Loop ChatRepair GPT-3.5 Defects4J v1.2/2.0 pass@500 162/337 single-function bugs
@@
Procedural Human-in-the-Loop CREF GPT-4 TutorCode avg@5 76.60% requires tutor guidance
@@
Procedural RAG-in-the-Loop Agentless GPT-4o SWE-bench Lite pass@1 32%
RAGFix Llama3 70B HumanEvalFix pass@1 78%
@@
Procedural AAG-in-the-Loop Repilot Incoder 7B Defects4J v1.2/2.0 pass@5000 116/273 single-hunk bugs
FormHe Gemma 2B Student Submissions pass@1 94.20%
PredicateFix Claude-3.5 Sonnet CVE-2020-13946 pass@1 81/117
@@
Agentic Tool-Augmented SWE-Agent GPT-4 Turbo SWE-bench Lite pass@1 18.00%
@@
Agentic LLM-as-Judges VulSifter GPT-4 CleanVul F1 82%
LLM4PatchCorrect StarCoder-7B Defects4J v1.2/2.0 AUC 76.9
Execution-free Critic Claude-3 Opus SWE-bench F1 91.60%
PatchDiff GPT-4o-mini SWE-bench-Verified accuracy 28.60%
Agentic Self-Controlled System MAGIS GPT-4 SWE-bench Lite pass@1 16.67%
@@
6. Open Challenges
Despite rapid progress, LLM-based APR still faces several open problems that cover the four paradigms.
Training and deployment cost. Fine-tuning large checkpoints needs many GPUs, long wall-clock hours, and sizable energy budgets. Even token-efficient adapters must still process millions of tokens. During inference, procedural and especially agentic workflows frequently call the LLM, invoke external tools, and rerun tests. Each round consumes thousands of tokens; running the loop locally demands high-memory GPUs, while sending it to a commercial LLM incurs substantial token charges. These long, costly sessions often exceed the few-minute latency and budget limits typical of continuous-integration pipelines, which makes real-world deployment challenging.
Evaluation reliability. Benchmarks often overlap with pre-training data, which causes data leakage and inflated scores. Weak oracles let a patch pass if any test succeeds, and underspecified test suites can miss semantic faults. This leads to patch overfitting: the model produces a plausible patch that passes every available test but still contains hidden errors. Stricter suites, extra hidden tests, and leakage-excluded datasets frequently reveal large accuracy drops. Some studies attempt to compensate by having experts manually judge patch correctness; however, manual review introduces reviewer bias and does not scale due to the cost of human verification.
Workflow robustness. The repair process is highly sensitive to multiple factors. Minor changes in prompt wording, sampling temperature, or random seeds can change the generated patch and may turn a successful repair into a failure. In agentic workflows, this randomness accumulates across steps, so two runs of the same system can follow different tool sequences and yield different results. Language models may also produce invalid commands, ignore provided rules, or loop too long, which makes the total running time unpredictable. These issues make it hard to guarantee repeatable outcomes or keep running costs within fixed limits in real-world pipelines.
Retrieval and context management. Most RAG pipelines still rely on embedding-based similarity search, which often misses some of the key files or functions and gives no recall guarantee. Even when relevant snippets are found, the added context can push the prompt beyond the LLM’s context limitation, limiting the effectiveness of RAG. Recent adaptive-memory systems, A-MEM, store each interaction as a structured note, maintain a link graph, and supply only a compact top-$k$ context, thereby improving recall while staying within the token budget (Xu et al., 2025). AAG pipelines generate diagnostics and traces directly from program analysis; however, current LLMs treat this structured data as plain text and fail to exploit its logical structure in a meaningful way.
Security and trust. LLM-generated patches often change more code than needed: they may pass the available tests yet hide latent bugs, break style guides, or slow the program. Calls to commercial APIs can leak proprietary code, and agents that bypass static checks may leave vulnerabilities unfixed. In addition, LLMs usually provide a natural-language rationale for each patch; if this explanation is incomplete or inconsistent with the actual code change, it can mislead reviewers and hinder debugging. A central open problem is how to verify both the patch and its rationale quickly and accurately, ideally before deployment, using lightweight static analysis, selective dynamic checks, and automated rollback when risks are detected.

### v2

@@ 7.3. Self-Controlled System
Overview Future research in agentic repair should emphasize cost-aware planning that caches intermediate results, prunes unproductive branches, and calls judgment modules only when their expected benefit justifies the extra computational resources, thereby reducing latency. Robust context management is also essential: agents must retain only the most relevant code fragments, tests, and analysis artifacts so that large repositories remain within the model’s context window without losing critical detail. Reliability will improve when lightweight static checks and policy guards intercept infinite loops or unsafe edits before they reach version control, satisfying continuous integration requirements. Finally, a tiered workflow that applies inexpensive heuristics to straightforward defects and escalates more complex cases to a fully autonomous agent can strike a balance between efficiency and autonomy. Sustained attention to these directions can elevate current prototypes into dependable tools for everyday software engineering practice.
Having surveyed fine-tuning, prompting, procedural pipelines, and agentic frameworks, we now synthesize how the control paradigms and augmentation layers jointly organize the design space and answer RQ2.
Figure 6. Alignment between LLM-based repair paradigms and typical deployment scenarios.
Answer to RQ2 The LLM-based software repair design space is organized into four control paradigms with retrieval and analysis augmentation as cross-cutting layers. Fine-tuning aligns a model to recurring repair patterns and works well for single-hunk or domain-specific defects. Prompting enables rapid fixes for single-function bugs without training, but is sensitive to prompt design and context limits. Procedural pipelines script generates and tests or retrieves and analyzes loops to deliver deterministic behavior on harder tasks. Agentic frameworks plan, call tools, and coordinate multi-step changes, which extend to multi-hunk and cross-file repository-level repairs. Figure 6 further maps these paradigms to representative deployment scenarios, including educational assignments, function-level benchmarks, repository-level issue suites, security vulnerabilities, and industrial CI workflows, indicating where each paradigm tends to be a primary or secondary fit. Mapping the representation works to these cells makes the control authority, iteration budget, and external-tool reliance explicit and ties each cell to the defect scopes it is best suited to handle.
8. Trends in Benchmark Repair Success Rates
This section consolidates published success rates for widely used software repair benchmarks to illustrate how performance evolves across paradigms. We report each system’s headline metric as initially published and record the evaluation assumptions alongside the score, rather than simply providing a single result. This maintains fidelity to the original evaluation setups and makes explicit differences in fault localization, scope, and pass@$k$.
We cover four widely-used benchmarks: Defects4J (v1.2 and v2.0), HumanEval-Java, SWE-bench Verified, and SWE-bench Lite. For each benchmark, we select representative systems that cover four paradigms and list their published scores along with the original metric and assumptions.
We use pass@k to denote the number of candidate patches for which at least one of k independent samples yields a patch that passes the test suites. When a wall-clock or token budget is enforced, we state it explicitly, for example, pass@10 within 5 hours.
Table 5. Defects4J trend snapshot.
Year Paradigm System Metric Notes
2022 Prompting (zero-shot) AlphaRepair top@$k$ within 5 hours: 74/109 Perfect fault localization.
2023 Prompting (few-shot) Xia et al. pass@200: 99/255 Perfect fault localization.
2023 Full fine-tuning Huang et al. pass@5: 82/107 Perfect fault localization.
2023 Full fine-tuning Jiang et al. pass@10: 31/138 Perfect fault localization.
2023 AAG-in-the-loop Repilot pass@5000: 116/273 Single-hunk Java bugs.
2023 Knowledge distillation KNOD pass@1000: 122/837 Perfect fault localization.
2024 AAG-enhanced prompting D4C pass@10: 180/437 Single-function setting, no FL hints.
2024 Test-in-the-loop ChatRepair pass@500: 162/337 Single-function setting.
2024 Tool-augmented agent RepairAgent pass@1: 164/835 Repository-level execution with tools.
2024 Distillation (non-AR student) NARRepair pass@1: 110/127 Perfect fault localization.
2025 LLM-as-Judges TSAPR pass@16: 201/835 Perfect fault localization.
On Defects4J (Table 5), the field evolves from early single-turn prompting with perfect fault localization toward analysis-augmented prompting, distillation, and procedural loops that reuse failing assertions or traces. Solved-bug counts rise substantially even as some assumptions are relaxed, for example moving from single-hunk Java snippets to single-function or repository-level settings.
Table 6. HumanEval-Java trend snapshot.
Year Paradigm System Metric Notes
2023 PEFT RepairLLaMA pass@10: 67.28% Perfect fault localization.
2024 PEFT Li et al. pass@10: 68.10% Perfect fault localization.
2024 PEFT (federated) Luo et al. pass@10: 76.88% Privacy-preserving training.
2025 PEFT (iterative) Ruiz et al. pass@10: 78.53% Perfect fault localization.
2025 PEFT (multi-objective) MORepair pass@10: 77.90% Rationale-aware training.
2024 Context-enriched FT NTR pass@100: 83.44% Template guidance, perfect FL.
2024 Test-in-the-loop ContrastRepair pass@40: 84.05% Contrastive test pairs.
2025 Analysis-augmented prompting TracePrompt accuracy: 71.30% Concise execution traces.
On HumanEval-Java (Table 6), parameter-efficient fine-tuning already delivers pass@10 near eighty percent. Context- and template-enriched fine-tuning, analysis-augmented prompts with concise traces, and conversational test-in-the-loop loops push headline scores higher at larger k. These results suggest that most low-hanging fruit on this benchmark now lies in better context selection and decoding strategies rather than in further scaling model size.
Table 7. SWE-bench trend snapshot. Verified and Lite variants are reported separately.
Year Variant Paradigm System Metric
2024 Lite Tool-augmented agent SWE-Agent pass@1: 18.00%
2024 Lite Tool-augmented agent OpenHands pass@1: 26.00%
2024 Lite Tool-augmented agent AutoCodeRover pass@1: 19.00%
2024 Lite RAG-in-the-loop Agentless pass@1: 32.00%
2024 Lite Self-controlled agent SWE-Search pass@1: 39.00%
2025 Lite RAG-in-the-loop KGCompass pass@1: 46.00%
2025 Lite LLM-as-Judges TSAPR pass@32: 54.67%
2025 Lite LLM-as-Judges SpecRover pass@1: 31.00%
2024 Verified Human-in-the-loop HULA pass@1: 31.00%
2025 Verified RLFT + procedural SWE-RL pass@1: 41.00%
On SWE-bench (Table 7), repository-level repair benefits most from richer control and context. Tool-augmented agents achieve the first gains on SWE-bench Lite, but RAG-in-the-loop pipelines and knowledge-graph-guided retrieval drive pass@1 into the mid-forties by supplying focused repository context between generations. On SWE-bench Verified, human-in-the-loop pipelines and reinforcement-learning-enhanced procedural workflows raise success rates. Overall, progress on SWE-bench is closely tied to better repository modeling and feedback rather than to single-step prompting.
Heterogeneous metrics and sampling budgets further obscure whether improvements arise from better models, from more generous pass@k, or from stronger evaluation setups. Table 7 makes protocol differences explicit, supports more transparent comparison across systems, and highlights where bias is most likely to appear when interpreting published repair success rates.
Answer to RQ3 Across the benchmarks we survey, reported repair rates are shaped as much by protocol design as by model quality. On Defects4J and HumanEval-Java, many systems evaluate on different subsets of bugs and mix pass@$k$ with accuracy-style metrics, and most headline numbers still assume perfect fault localization rather than end-to-end repair. On SWE-bench, recent agentic and reinforcement-learning-enhanced systems are converging on repository-level, end-to-end settings with fixed pass@1 budgets, but vulnerability repair remains fragmented across small, non-overlapping datasets, so scores there are substantially harder to compare. These observations suggest three concrete priorities for future evaluation. First, results should be reported on the full version of each benchmark, or with clearly marked subsets and explicit denominators to make the results comparable. Second, studies should standardize metrics, for example, always reporting pass@1 and one shared larger $k$ without other sampling or wall-clock budget, and leveraging end-to-end settings without perfect fault-localization hints. Third, the vulnerability-repair community would benefit from a consolidated benchmark in the spirit of SWE-bench, with curated repository-level defects and strong oracles; until such a suite exists, improvements on narrow vulnerability datasets should be treated as evidence of local progress rather than as general advances in LLM-based repair.
9. Related Work
Earlier surveys on LLM-based software repair focus on predate LLM-based repair or mention LLMs only briefly, so they do not analyze LLM-centered control structures. Huang et al. (Huang et al., 2023b) updated the traditional landscape with a four-way taxonomy over search, constraint, template, and learning techniques, but relegated LLMs to a short remark about future potential. Zhang et al. (Zhang et al., 2023a) presented a thorough synthesis of learning-based software repair, framing the workflow around localization, generation, ranking, and validation; because their literature search concluded before mid-2022, large-scale LLMs such as Codex receive only brief mention, and prompt-driven repair workflows are not yet covered in depth. Winter et al. (Winter et al., 2022) shifted focus to developer perceptions and usability, without a technical taxonomy. Dikici et al. (Dikici and Bilgin, 2025) grouped tools into template, machine-learning, and deep-learning categories, implicitly folding GPT-style systems into the deep-learning bin without analyzing their distinctive traits or their impact on the repair loop.
More recent work concentrates on LLMs and related evaluation issues. Anand et al. (Anand et al., 2024) explicitly surveyed LLM-based software repair and code generation but covered only a small set of software repair papers and did not propose a dedicated taxonomy. Haque et al. (Haque et al., 2025a) conducted a systematic literature review on parameter-efficient fine-tuning for LLMs, covering software repair but focusing on fine-tuning methods rather than repair workflows. He et al. (He et al., 2025) examined the emerging LLM-as-judge paradigm for code evaluation, offering insights into assessing LLM-generated patches while omitting a repair-oriented taxonomy. Zhou et al. (Zhou et al., 2024a) published a review that catalogs work on vulnerability detection and repair and discusses adaptation techniques and dataset usage, but its analysis is limited to security-related bugs, and it treats retrieval-augmented and agentic methods mainly as future directions. Zhang et al. conducted a literature review focused on LLM-based software repair (Zhang et al., 2024a), collecting a broader corpus of LLM-based repair studies and classifying them by LLM utilization modes, without explicitly modeling control authority or augmentation layers. Zhang et al. provide a critical review that uses ChatGPT as a case study on a contamination-resistant benchmark, examining prompt design and data-contamination risks (Zhang et al., 2023b). Table 8 summarizes these surveys and their coverage of LLM-based techniques.
Table 8. Recent surveys on automated program repair and their coverage of LLM-based techniques.
Survey Venue Year LLM Coverage Classification Scheme Key Limitations
Winter et al. (Winter et al., 2022) IEEE TSE 2022 None Human-factor perspective Lacks technical classification; predates LLM era.
Huang et al. (Huang et al., 2023b) arXiv 2023 Mentioned as future work Search, constraint, template, learning No category devoted to LLM repair.
Zhang et al. (Zhang et al., 2023a) ACM TOSEM 2023 Moderate Pipeline stages Excludes non-ML approaches and prompt-based repair.
Zhang et al. (Zhang et al., 2023b) arXiv 2023 Strong for ChatGPT case ChatGPT-centered critical review and prompt variants Single-model case study; not a general software repair taxonomy or broad LLM-based software repair survey.
Anand et al. (Anand et al., 2024) arXiv 2024 Strong LLM-based software repair and code generation software repair coverage limited; no fine-grained taxonomy.
Zhang et al. (Zhang et al., 2024a) arXiv 2024 Strong LLM utilization modes and repair scenarios Emphasizes how LLMs are used rather than control paradigms; does not explicitly model retrieval- or analysis-augmented repair; limited cell-wise comparison of evaluation practices.
Dikici et al. (Dikici and Bilgin, 2025) KIS 2025 Partial Template, ML, deep learning LLM tools subsumed under deep learning.
Haque et al. (Haque et al., 2025a) arXiv 2025 Moderate Fine-tuning strategy taxonomy Focused on fine-tuning methods, not repair workflows.
He et al. (He et al., 2025) arXiv 2025 Evaluation only Evaluation role taxonomy Covers judging, not generation or workflow design.
Zhou et al. (Zhou et al., 2024a) ACM TOSEM 2025 Strong Security-oriented vulnerability repair Focuses only on vulnerabilities; no agentic approaches
Across these studies, three limitations remain. First, most taxonomies either predate LLM-based repair or retain a pipeline- or model-centric view and therefore do not treat prompt-driven, retrieval-augmented, procedural, and agentic workflows as first-class control paradigms tied to who steers the repair loop and whether the base LLM is adapted. Second, LLM-based works are often omitted or folded into broad deep-learning bins, which obscures how foundation models reshape software repair and how external tools, retrieval modules, and analysis components are orchestrated around them. Third, evaluation challenges specific to LLMs, including data leakage, weak test suites, heterogeneous pass@k settings, and differing fault-localization assumptions, are rarely treated systematically or aligned across benchmarks. We complement these surveys by offering a control-paradigm-centered taxonomy defined by parameter adaptation and control authority with explicit retrieval- and analysis-augmentation layers, by mapping 62 recent systems into this design space together with their defect scope and deployment scenario, and by consolidating benchmark, metric, and fault-localization assumptions into comparative tables, supported by a versioned, auditable artifact that exposes screening and coding decisions for future reuse.
10. Open Challenges
Despite rapid progress, LLM-based software repair still faces several open problems that span the four paradigms.
Training and deployment cost. Fine-tuning larger LLMs requires many GPUs and long wall-clock hours. Even token-efficient adapters still process millions of tokens. During inference, procedural and especially agentic workflows frequently call the LLM, invoke external tools, and rerun tests. Each round consumes a significant amount of tokens; running the loop locally demands high-memory GPUs, while sending it to a commercial API incurs substantial token charges. These long, costly sessions make real-world deployment of repair systems challenging.
Evaluation reliability. Benchmarks may overlap with pretraining data, which causes data leakage and inflated success rates. Weak test suites let a patch pass if all test succeeds, and underspecified test suites can miss semantic faults of LLM-generated patches. This leads to patch overfitting: the LLM produces a plausible patch that passes every available test but still contains hidden defects. Stricter suites, extra hidden tests, and leakage-excluded datasets frequently reveal accuracy drops. Some studies attempt to compensate by having experts manually judge patch correctness; however, manual review introduces reviewer bias and does not scale due to the cost of human verification.
Workflow robustness. Minor changes to prompt wording, sampling temperature, or random seeds can alter the generated patch and turn a successful repair into a failure. In agentic workflows, this randomness accumulates across steps, so two runs of the same system can follow different tool sequences and yield different results. Language models may also produce invalid commands, ignore provided rules, or loop for too long, making the total running time unpredictable. These issues make it hard to guarantee repeatable outcomes or keep running costs within fixed limits in real-world pipelines.
Retrieval and context management. Most RAG pipelines still rely on embedding-based similarity search, which often misses some of the key files or functions and gives no recall guarantee. Even when relevant snippets are found, the added context can push the prompt beyond the LLM’s context limitation, limiting the effectiveness of RAG. Recent adaptive memory systems, such as A-MEM, store each interaction as a structured note, maintain a link graph, and supply only a compact top-$k$ context, thereby improving recall while staying within the token budget (Xu et al., 2025b). AAG pipelines generate diagnostics and traces directly from program analysis; however, current LLMs treat this structured data as plain text and fail to meaningfully exploit its logical structure.
Security and trust. LLM-generated patches often change more code than needed: they may pass the available tests yet involve new latent bugs, break style guides, or slow the program. Calls to commercial APIs can leak proprietary codes and documents, and agents that bypass static checks may leave vulnerabilities unfixed. In addition, LLMs usually provide a natural-language rationale for each patch; if this explanation is incomplete or inconsistent with the actual code change, it can mislead reviewers and hinder debugging. A central open problem is how to verify both the patch and its rationale quickly and accurately, ideally before deployment, using lightweight static analysis, selective dynamic checks, and automated rollback when risks are detected.
Answer to RQ4 Across paradigms, four bottlenecks stand out for semantic correctness beyond tests and for repository-level multi-hunk repair: (i) limited and often shallow test suites, (ii) brittle and expensive multi-step workflows whose time and token usage grow quickly with project size, (iii) retrieval and analysis pipelines that still supply noisy or incomplete context, and (iv) weak assurance mechanisms for the safety and trustworthiness of accepted patches. We see three concrete opportunities. First, build benchmarks and test suites that explicitly target large-scale, multi-hunk changes, with stronger oracles, hidden tests, and contamination-aware splits, so that progress reflects real semantic correctness rather than overfitting narrow unit suites. Second, design cost-aware repair policies that treat success per unit budget as an objective, by reusing intermediate analysis results, capping the number of LLM calls, and falling back to simpler scripted pipelines when a bug appears easy, thereby controlling latency and variability on large repositories. Third, separate generation from evaluation more sharply by combining analysis- and test-based modules that gate patch acceptance and rationales, so that final decisions do not rely solely on larger LLMs.
Table 9. Detailed Information of LLM-based Software Repair Systems
@@ 10. Open Challenges
Fine-Tuning
Full Fine-Tuning
Huang et al. Recoder dataset CodeT5 Defects4J v1.2/2.0 pass@5 82/107 Perfect Fault Localization
Jiang et al. Recoder dataset Incoder 7B Defects4J v1.2/2.0 pass@10 31/138 Perfect Fault Localization
@@
Fine-Tuning
PEFT
RepairLLaMA Megadiff CodeLlama 7B HumanEval-Java pass@10 67.28% Perfect Fault Localization
@@
Luo et al. TutorCode CodeLlama 13B HumanEval-Java pass@10 76.88%
Li et al. APR-Instruction DeepSeek-Coder 7B HumanEval-Java pass@10 68.10% Perfect Fault Localization
Ruiz et al. Recoder dataset DeepSeek-Coder 7B HumanEval-Java pass@10 78.53% Perfect Fault Localization
Fine-Tuning
Knowledge Distillation
KNOD GitHub / Defects4J v1.2/2.0 pass@1000 122/837 Perfect Fault Localization
DistiLRR MBPP CodeLlama-7B HumanEval-Perl pass@1 42.10% train/test split
@@
Fine-Tuning
RLFT
RePair CodeNet4Repair StarCoder-16B CodeNet4Repair pass@5 65.66% train/test split
@@
Vul-R2 PrimeVul Qwen-14B-Instruct-1M PrimeVul pass@10 24.83 train/test split
Fine-Tuning
Context-Enriched
TraceFixer CodeNet CodeT5-small CodeNet pass@10 87%
InferFix InferredBugs Codex-12B InferredBugs pass@1 76.80% train/test split
@@
Prompting
Zero-Shot
AlphaRepair CodeBERT Defects4J v1.2/2.0 top@k (5h) 74/109 Perfect Fault Localization
Prenner et al. Codex-175B Quixbugs pass@1 37/80 Perfect Fault Localization
Fan et al. Codex-175B LMDefects pass@5 46/113
Tian et al. GPT-3.5 Turbo LeetCode pass@5 105/120
Prompting
Few-Shot
Xia et al. Codex 12B Defects4J v1.2 pass@200 99/255 Perfect Fault Localization
Gao et al. Codex-175B B2F small full match 32.25%
Ahmed et al. Codex-175B MODIT pass@30 31.80%
@@
Prompting
+ RAG
Ehsani et al. Llama-3.3 BugsInPy pass@5 79.62% Perfect Fault Localization
@@
Prompting
+ AAG
D4C GPT-4 Defects4J v1.2/2.0 pass@10 180/437 single-function bugs
@@
Procedural
Test-in-the-Loop
ChatRepair GPT-3.5 Defects4J v1.2/2.0 pass@500 162/337 single-function bugs
@@
Procedural
Human-in-the-Loop
CREF GPT-4 TutorCode avg@5 76.60% requires tutor guidance
@@
Procedural
RAG-in-the-Loop
Agentless GPT-4o SWE-bench Lite pass@1 32%
PATCH GPT-4 BFP pass@5 39.81%
@@
Procedural
AAG-in-the-Loop
Repilot Incoder 7B Defects4J v1.2/2.0 pass@5000 116/273 single-hunk bugs
SAN2PATCH GPT-4o VulnLoc pass@5 79.49%
PredicateFix Claude-3.5 Sonnet CVE pass@1 81/117
@@
Agentic
Tool-Augmented
SWE-Agent GPT-4 Turbo SWE-bench Lite pass@1 18.00%
@@
Agentic
LLM-as-Judges
Abstain and Validate Gemini 2.0 Flash Google accept@1 62%
TSAPR GPT-3.5 Defects4J v1.2/2.0 pass@16 24.07%
SpecRover Claude-3.5 Sonnet SWE-bench Lite pass@1 31.00%
Agentic
Self-Controlled System
MAGIS GPT-4 SWE-bench Lite pass@1 16.67%
@@
11. Threats to Validity
11.1. Internal validity
Study identification, selection, and coding threaten internal validity; we mitigate these threats with a structured pipeline. We search multiple digital libraries with calibrated queries and citation chasing, apply explicit inclusion and exclusion criteria, and use a unified filtering process reviewed by a coauthor with extensive program repair experience to ensure consistent treatment of borderline cases. All search scripts and results are released in our replication package, enabling others to inspect our decisions and recompute statistics with the same notebooks.
11.2. External validity
Narrow corpora and artificial evaluation settings threaten external validity. Representation intentionally focuses on work published in tier-1 SE, AI, and security venues, and on studies that report quantitative results on large-scale benchmarks. As a result, tools evaluated only on proprietary data fall outside our scope. We report headline scores as published initially and do not normalize pass@k or budgets across studies. We also do not perform a joint reevaluation. This preserves faithfulness to prior work but limits cross-paper comparability. To mitigate this limitation, we annotate benchmark scope, metric definitions, oracle strength, and fault localization assumptions alongside each reported score to reduce misleading.
11.3. Construct validity
Ambiguous definitions and measurements threaten construct validity. We mitigate these issues by operationalizing constructs precisely and enforcing a stable coding scheme. We define the four repair paradigms and cross-cutting augmentation layers by workflow control and accessible information, apply them consistently across all studies, and discuss borderline cases to reach a shared decision.